Part 1 -

The Tucker family has health insurance coverage that pays 65 percent of out-of-hospital expenses after a deductible of $600 per person. If one family member has doctor and prescription medication expenses of $2,100, what amount would the insurance company pay?

Part 2 -

A health insurance policy pays 70 percent of physical therapy costs after a deductible of $380. In contrast, an HMO charges $24 per visit for physical therapy. How much would a person save with the HMO if he or she had 12 physical therapy sessions costing $95 each?
Answer #1

Part 1:

doctor and prescription medication expense = $2,100

there is a deductible of $600 per person and after the deduction the insurance pays 65%

We will deduct $600 from $2100 and calculate 65% of the difference

This will be = (2100 -600 ) * 65% = (1500) * 65% = $975

The insurance should pay $975

Part 2:

Firs let us calculate the total cost of 12 therapy seasons = 12 * $95 = $1,140

Option 1 - insurance pays 70% after deductible of $380. This means we have to subtract $380 from the total cost and then calculate 70%

insurance pays = (1140-380) * 70% = $532

Hence the patient pays 1,140 - 532 = $608

Option 2: HMO charges $24 per visit. Hence total amount the patient pays = 12 *$24 = $288

The person will save $608 - $288 = $320